OpenStudy (yourboydusty):

What is the distance between the two points graphed below? If necessary, round your answer to two decimal places.

OpenStudy (yourboydusty):

@YanaSidlinskiy

OpenStudy (yanasidlinskiy):

To find the length of a, you calculate the difference between the y values. In other words, you go from -4 to -2, which is a distance of *how many units?* For b, it's the same thing except with x values. You go from 1 to 3 which is how many units?

OpenStudy (yourboydusty):

umm...3 units

OpenStudy (yanasidlinskiy):

Once you find the difference between those, you need to square them, using: a^2 + b^2 = c^2 (where c is the hypotenuse) a^2 + b^2 = (distance between points)^2

OpenStudy (yanasidlinskiy):

There's 3 units and what's the other one?

OpenStudy (yourboydusty):

2?

OpenStudy (yanasidlinskiy):

3^2 is 9 2^2 is 4 4+9 is 13.. Square root of 13 is about 3.61

OpenStudy (yourboydusty):

So that is it?

OpenStudy (yanasidlinskiy):

@Elsa213 @ShadowLegendX Is that right? Yes I think so

OpenStudy (shadowlegendx):

@YanaSidlinskiy Are you trying to solve for this, like it's a right triangle Using Pythagorean's Theorem?

OpenStudy (yanasidlinskiy):

Yea... Don't tell me I'm wronggggg

OpenStudy (shadowlegendx):

I won't, but I'll do it correctly

OpenStudy (yanasidlinskiy):

So it's wrong?

OpenStudy (yourboydusty):

thank you for trying Yana

OpenStudy (yanasidlinskiy):

I mean aren't you supposed to use and draw it in like a right triangle for it to be easier?

OpenStudy (shadowlegendx):

(-2,-4) (3, 1) (x, y) The difference between both the x, and the y, is 5 units. a and b, in a^2 + b^2 = c^2, would both be 5.

OpenStudy (yanasidlinskiy):

So that's where you square them both? And then finish it off like I did?

OpenStudy (shadowlegendx):

\[a^2 + b^2 = c^2\] \[5^2 + 5^2 = c^2\] \[25 + 25 = c^2\] \[50 = c^2\] Take the root, of both sides \[c = \sqrt 50 \rightarrow c = 7.07106781187\]

OpenStudy (yourboydusty):

oh man im getting a headache....i so dont like math lol

OpenStudy (shadowlegendx):

What part do you not understand?

OpenStudy (yourboydusty):

just the last part

OpenStudy (shadowlegendx):

http://prntscr.com/ctu8wk

OpenStudy (yanasidlinskiy):

Hey it's Pythagrean Theorem Or however you spell it lol.. THat's what we just did except with different numbers....

OpenStudy (shadowlegendx):

We used Pythagorean Theorem, which is \[a^2 + b^2 = c^2\] Where, a and b are the two sides, and c is the hypotenuse, or the sloped line that we are solving the distance for

OpenStudy (yourboydusty):

ugh and i am no were near finishing my work i have over 50 questions and i only have 24 done ;_;

OpenStudy (yourboydusty):

okay... i get that

OpenStudy (shadowlegendx):

We just found the two sides by observing the coordinates, then solved for the third side

OpenStudy (yourboydusty):

So the answer is.....

OpenStudy (yanasidlinskiy):

7.18 .....

OpenStudy (shadowlegendx):

If you understood this, you would know I gave it to you 5min ago ;_;

OpenStudy (yanasidlinskiy):

Hey I understood Ig? Lol

OpenStudy (shadowlegendx):

c = 7.07

OpenStudy (shadowlegendx):

rounded to two decimal points

OpenStudy (yourboydusty):

oh....srry im a little slow on the uptake :P